Asinger finds herself with extra $5000 at the end of the year. Calculate the annual rates of return for each of the following investments (rounding to the nearest 0.1% where needed).

a) she can renovate her home to make it more energy-efficient. Doing so will save her on average $100 per month on her energy bills

Rate of return:

b) she has a $12,000 of student loan debt, on which she is charge an annual yield of 6.8%. She could reduce this debt to $7,000

Rate of return:

c) She can loan the money to a relative's business, who promises to pay her back $10,000 in 6 years (Assume this loan is safe - she's 100% certain that she'll be paid back.)

Rate of return:
Answer #1

a) Monthly return = (100/5000) = 2%

Annual return = (1.02)^12 -1 = 26.8%

b) Since the singer will save annual yield is 6.8%, her rate of annual return is 6.8%

c) Rate of return = (10000/5000)^(1/6) -1 = 12.24%